# Towards Swift Serverless LLM Cold Starts with ParaServe

### v1

@@ 4 Cluster-Level Controller
In this section, we present the design of the cluster-level central controller in ParaServe. The controller makes decision on two aspects: (a) pipeline parallelism size and (b) worker placement. To find the optimal pipeline parallelism size, we first analyze the tradeoffs involved in pipeline parallelism, and then develop an algorithm to select the appropriate parallelism size based on the model size and user SLOs (§4.1). For worker placement, we employ a network-contention-aware worker placement strategy to prevent SLO violations (§4.2).
4.1 Parallelism Size Selection
@@
Figure 4(a) illustrates the normal cold-start workflow, which goes through six stages sequentially. As model fetching is the most time-consuming stage, ParaServe leverages pipeline parallelism to optimize cold starts. As shown in Figure 4(b), upon a cold start, we launch multiple workers on different GPU servers, each fetches a part of the model. During inference, workers exchange intermediate results but the size of these messages is relatively small. By reducing the portion of the model each worker needs to fetch, pipeline parallelism significantly reduces cold start latency.
4.1.1 Tradeoff Analysis
@@
Despite the benefits of pipeline parallelism, it introduces resource and performance overheads. To better understand these tradeoffs, we conduct experiments using a testbed with four GPU servers, each equipped with a NVIDIA A10 GPU and 188GB memory. The network bandwidth of each server is 16Gbps.
@@
Figure 5(a) demonstrates the TTFT of models under different pipeline parallelism sizes. Larger parallelism sizes reduce model fetching time, resulting in shorter TTFT. Note that the marginal improvement on TTFT diminishes as other cold-start stages also consume non-negligible time. This limitation is further tackled by worker-level optimizations in §5.
In terms of TPOT, pipeline parallelism has a negative impact as shown in Figure 5(b). This is because individual workers in a pipeline parallelism group allocate GPU memory proportionally to their partitioned model size by default, and the amount of allocated memory also determines the GPU computation quota in our production environment. Thus, reduced GPU memory leads to severe performance degradation.
Fortunately, users can allocate more GPU resources to workers than the default quota to achieve lower TPOT. Figure 5(c) illustrates the inference performance when different amount of GPU memory is allocated to each worker and pipeline parallelism size is four. Note that the performance degradation would be negligible if each worker in parallelism group is given the same amount of GPU memory as in the non-parallelized case (i.e., parallelism size is one).
@@
In conclusion, larger parallelism size reduces the cold start latency, but impacts the inference performance. Allocating more GPU memory to workers reduces the performance loss but at the cost of higher price. ParaServe designs the pipeline parallelism size selection algorithm based on these two tradeoffs.
Input: time cost of container creation and runtime initialization , data transmission , prefill , decoding , and waiting ; model size ; GPU server network bandwidth and PCIe bandwidth ; user specified requirements and .
Output: cold start scheme including pipeline parallelism size , #full-memory workers , and selected GPU servers .
4.1.2 Algorithm Design
The parallelism size selection algorithm takes model size, user SLOs, and cluster resource status as inputs. The cluster resource status includes the remained GPU memory of all servers and their network bandwidth, adjusted by network-contention-aware worker placement policy in §4.2. The algorithm output is a cold start scheme containing the pipeline parallelism size, GPU memory usage of workers, and worker placement.
@@ 4.1.2 Algorithm Design
To narrow down the search space, we limit the possible GPU memory usage of a worker to two cases: (a) the same as the non-parallelized setup (b) proportional to the inverse of pipeline parallelism size. We call them full-memory workers and low-memory workers, respectively.
@@
Our algorithm has two stages. First, we enumerate all choices with different pipeline parallelism sizes or GPU memory usage and predict the TTFT and TPOT for each choice. Second, we select the optimal choice that satisfies user SLOs with minimum resource usage. The maximum pipeline parallelism size is limited to as larger parallelism sizes gain little improvement.
@@
The TTFT and TPOT prediction takes historical information as inputs, including the time cost of container creation and runtime initialization (), data transmission (), prefill (), and decoding (). The data transmission time is the latency of TCP network between GPU servers. The prefill and decoding time cost are model-specific metrics obtained from the model’s previous executions. The maximum waiting time of requests for the model is also needed, defined as . Assume the pipeline parallelism size is and there are full-memory workers , while the network and PCIe bandwidth of servers that workers reside in are and . Finally, we predict the TTFT using the following equation:
@@
where is the model size. TTFT is made up of five parts: request waiting, runtime initialization, model fetching, model loading, and prefill. The model fetching time is influenced by the partitioned model size and the slowest worker, while the prefill time depends on parallelism size and number of full-memory workers. A full-memory worker costs units of time in prefill, while a low-memory worker costs units of time.
@@
To satisfy TTFT requirement, we select GPU servers to minimize the model fetching and loading time for a given pipeline parallelism size and number of full-memory workers , Assume the GPU servers that can accommodate full-memory workers are . Apart from them, there are also servers that can accommodate low-memory workers, denoted as . Our selection strategy is to first allocate the top servers with minimum model fetching and loading time (i.e., the smallest ) from to full-memory workers, and then merge the remaining servers into the set . After that, we select the top servers from the merged set by same strategy and allocate them to all low-memory workers.
Similarly, we can predict the TPOT as follows:
Among all choices that satisfy both TTFT and TPOT SLOs, ParaServe selects the one with minimal parallelism size and GPU resource usage. Algorithm 1 presents the pipeline parallelism selection algorithm.
@@ 4.2 Network-Contention-Aware Worker Placement
The second type of contention is the interference between cold-start workers on the same GPU server. Cold-start workers that collocated on a GPU server will compete for network resources during model fetching, leading to unexpected cold start performance. In contrast, simply assigning cold-start workers to servers in a mutually exclusive manner results in low GPU resource utilization. To this end, we propose a network-contention-aware worker placement policy that identifies network contentions and places cold-start workers based on user SLOs.
4.2.1 Policy Design
Input: each worker’s deadline given by TTFT and estimated pending model size
Initially, collocated workers share the network bandwidth with equal credits. Once we assign a worker to a GPU server, we record the model size it needs to fetch and the user-specified maximum TTFT. To place a new worker, we inspect each GPU server to check that whether adding a cold-start worker to this server would lead to SLO violations for existing workers. If check passed, we estimate and record the network bandwidth that the new worker can obtain. Note that PCIe bandwith is much higher than network and PCIe switch is able to further isolate PCIe usage across tasks, so we do not take PCIe contention into account.
Specifically, we denote the bandwidth of a GPU server as . For each cold-start worker on the GPU server, we record the fetching deadline and estimate its pending model size . The fetching deadline comes from the prediction of TTFT (Eq. 1). When a new cold-start worker comes, the network bandwidth would change to if there were cold-start workers on the server. We check whether all cold-start workers on the server are able to complete fetching before deadline under the new bandwidth. Formally, we check that whether the following condition is true for all workers:
where represents current time. This server accepts the new worker if all workers passed the check.
To estimate the pending model size, we record the time of last network bandwidth change ’. The start and completion of a cold start change the network bandwidth, informing controller to adjust the pending model size. Here, we assume the current time is and the number of cold-start workers before change is . Then we adjust the pending model size upon each bandwidth change according to the following equation:
After adjustment, a that is below zero means that the worker has fetched the model ideally, thus we delete it from the cold-start worker list. Algorithm 2 outlines the pseudocode of obtaining server bandwidth and handling bandwidth change.
@@ 5 Worker-Level Overlapping
As described in §4.1.1, the TTFT of pipeline parallelism has diminishing marginal returns due to container initialization stages, including container creation, library loading and CUDA context initialization. Container creation is slow because function images of LLM workloads usually contain multiple Python libraries such as PyTorch [42], TensorFlow [48], TensorRT [46] and vLLM [44]. These Python libraries have complex dependencies, leading to large image sizes. For example, the basic image for running vLLM is around 8.31GB. Loading and initializing numerous libraries also cost long libray loading time. Preparing a universal container for all workers [49, 50, 51] does not help in this case because users have requirements for different library versions.
In this section, we present the worker-level overlapping in ParaServe that carefully reorganizes the startup workflow to reduce worker initialization time. Basically, ParaServe leverages two optimizations to overlap initialization stages. First, we prefetch model weights on local GPU server as soon as container creation starts so that the container initialization stages are overlapped with model fetching. Second, we prioritize CUDA context initialization and develop a parameter manager to load model while initializing Python libraries so that the library loading stage is overlapped with model loading.
@@ 5.1 Model Prefetching
ParaServe launches a model prefetcher on each GPU server, which is responsible for proactively fetching models from remote storage. When a worker has been allocated to the server, the central controller informs the model prefetcher about model metadata. After that, the prefetcher starts to load the model weights from remote storage to a shared memory region. The cold-start worker will fetch parameters from shared memory in a streaming manner after runtime has initialized.
Model prefetching starts before container creation, so that the container creation and runtime initialization stages are overlapped. The worker performs model prefetching and loading in a pipelined fashion. In the shared memory region of a model, we use the first eight bytes to store the address that represents the end of currently fetched model weights. Model weights are represented using SafeTensors format [52]. SafeTensors format contains the metadata of all parameters at the beginning of the file, so that it is convenient for the worker to check whether a tensor has been fetched.
@@ 5.2 Parameter Manager
ParaServe leverages a parameter manager to load model parameters in the background. The parameter manager runs in an individual thread and is responsible for resolving tensor metadata, reading weights from the shared memory, and finally loading weights into GPU. The whole procedure is zero-copy and pipelined. The parameter manager also takes advantage of the high parallelism of GPU cores and uses multiple CUDA streams to load models. The priorities of CUDA streams are determined according to whether the loading process is on the critical path or in the background.
@@
Figure 6(a) illustrates the optimized cold-start workflow after adopting model prefetching and parameter manager. Model fetching and container initialization runs in parallel, while library loading is overlapped with model loading. In this way, the overhead of other cold start stages is concealed behind model fetching, reinforcing the effectiveness of pipeline parallelism.
As ParaServe may launch workers in pipeline parallelism groups and later load the remaining part of model in background, Figure 6(b) also shows the cold-start workflow in this scenario. The model prefetcher downloads two parts of model sequentially. After the first part of model has been loaded, the worker starts to perform inference with other workers in the pipeline parallelism group. Meanwhile, the parameter manager loads the second part of model in background and finally transform itself into a standalone worker.
TTFT prediction with worker-level overlapping. After applying worker-level optimizations, the TTFT is reduced by fine-grained overlapping. Based on additional historical information including the time cost of container creation (), CUDA context initialization () and library loading (), we change the TTFT prediction used in pipeline parallelism size selection (Eq. 1) into the following formula.
6 Pipeline Consolidation
ParaServe proposes pipeline consolidation that merges pipeline parallelism groups into individual workers to obtain optimal performance. Specifically, after the parameter manager has loaded the requested model parameters in pipeline-parallel inference, we inform it to continue loading the remaining part of model in background. Once loading completes, the worker returns to the non-parallelized setup and serves subsequent requests with whole model in local GPU, thereby achieving optimal performance. Note that the parameter manager loads the second part of model in low-priority CUDA streams, so that the performance of inference task will not be affected.
@@ 6.1 Worker Scaling
ParaServe provides two scaling choices. First, after workers have loaded corresponding parts of models, we can perform scaling down by allowing one of them to fetch the unloaded model parts in background. Once all parameters are loaded, we migrate all existing requests to that worker by migrating their key-value cache. Finally, the worker continues to generate tokens with whole model while other workers are terminated. Figure 4(c) demonstrates the process of scaling down. In this way, we obtain the first token earlier by parallelized model fetching and finally obtain a worker with whole model, similar to normal cold starts.
The second scaling choice is scaling up, changing all cold-start workers into individual serving endpoints, as Figure 4(d) has shown. LLM inference workloads usually have bursty traffic [27, 38, 53]. In serverless LLM serving, bursty requests often cause the system to concurrently create multiple workers in the cluster. In this scenario, the scaling up method allows for quickly creating workers in pipeline parallelism groups, so that the system can reach the maximum throughput earlier.
Figure 7 uses an example of creating new workers concurrently to show the benefits of scaling up. When a burst of requests arrive, existing serverless LLM serving systems initialize new workers concurrently, resulting in significant network and memory consumption in the cluster. Additionally, workers are not able to serve requests until the whole model has loaded, as shown in Figure 7(a). In contrast, ParaServe with scaling up is able to prepare pipeline parallelism groups quickly. Figure 7(b) demonstrates the startup strategy of ParaServe, which allocates workers to pipeline parallelism groups, where one group has two workers and others have four. In this way, workers are able to generate tokens upon loaded part of the model, thereby starting to serve requests much earlier. Since pipeline parallelism has comparable throughput with same number of workers that hold the whole model [38], the system can achieve the maximum possible throughput upon the creation of all groups. Furthermore, workers continue to load models in background, and eventually turn into individual workers to reduce inference latency. In other words, the system is able to achieve high throughput earlier and finally obtains same number of workers with normal cold starts.
In default, ParaServe adopts the scaling down mechanism to reduce cold start latency with little overhead. To quickly react to bursty loads and change to the scaling up mechanism, we determine the number of workers to remain with a sliding window strategy. For each model, we record the number of requests that arrived during the past window and use it as the predicted maximum number of requests that may arrive in the next window. The required number of workers to retain is calculated based on the current number of requests in the waiting queue combined with the predicted maximum number of requests expected to arrive in the next window. In cold start, we create a pipeline parallelism group that is no smaller than required number of new workers and later transform the group into desired number of individual workers. We may create multiple pipeline parallelism groups to serve a sudden burst of requests.

### v2

@@ 4 Cluster-Level Controller
In this section, we present the design of the cluster-level central controller in HydraServe. Figure 4(a) illustrates the normal cold-start workflow, which goes through six stages sequentially. HydraServe leverages pipeline parallelism to optimize model fetching, the most time-consuming stage. As shown in Figure 4(b), upon a cold start, we launch multiple workers on different servers, each fetching a part of the model. During inference, workers exchange intermediate results but the sizes of these messages are relatively small. By reducing the portion of the model each worker needs to fetch, pipeline parallelism significantly reduces cold start latency. HydraServe finds the optimal resource allocation scheme for each cold-start model based on the model size and user SLOs (§4.1), and employs a network-contention-aware worker placement strategy to prevent SLO violations (§4.2).
4.1 Resource Allocation
To design an efficient resource allocation algorithm for pipeline workers, we begin by quantifying the tradeoffs involved in pipeline parallelism. Based on this analysis, we propose a TTFT and worst-case TPOT prediction method.
@@ 4.1 Resource Allocation
Tradeoff analysis. The benefits of pipeline parallelism come with performance and resource overhead. To better understand these tradeoffs, we conduct experiments using four GPU servers, each equipped with a NVIDIA A10 GPU and 188 GB memory. The network bandwidth of each server is 16 Gbps.
@@
Figure 5(a) demonstrates the TTFT of models under different pipeline parallelism sizes. Larger parallelism sizes reduce model fetching time, resulting in shorter TTFT. Note that the marginal improvement in TTFT diminishes as other cold-start stages also consume non-negligible time. This limitation is further addressed by worker-level optimizations in §5.
Pipeline parallelism has a modest impact on TPOT, as shown in Figure 5(b). This is because the messages transmitted between workers are relatively small. For example, Llama2-7B incurs only 8 KB of inter-layer results per token.
However, pipeline parallelism may lead to worker colocations across different models. For a model initially requiring GPU memory, distributing it across GPUs reduces per-GPU memory consumption to . Under heavy load, the cluster must co-place workers whose combined memory requirements fit within a single GPU’s capacity to maximize resource utilization. This colocation of workers from different models can degrade performance.
Figure 5(c) illustrates this issue by changing the GPU memory allocated to each model, with pipeline parallelism size fixed at 4. Allocated GPU memory represents the cost of a model, and lower memory usage per model leads to worker colocation. For example, allocating 32GB GPU memory to each model makes three models share four GPUs. As the cost decreases, each GPU hosts multiple concurrently running workers. Consequently, the GPU’s computational resources are allocated proportionally to each worker’s reserved memory, resulting in longer TPOT.
@@
In conclusion, while increasing pipeline parallelism size reduces the cold start latency, it may impact inference performance due to GPU sharing. Allocating additional GPU memory to workers can mitigate this performance loss, but at the expense of higher resource consumption. HydraServe’s resource allocation algorithm systematically balances these tradeoffs to optimize system performance.
Algorithm design. The resource allocation algorithm finds an appropriate resource allocation strategy for each cold-start model according to user SLOs. Apart from pipeline parallelism size, the algorithm should decide how to place pipeline workers and allocate how much GPU memory to each worker.
@@
To narrow down the search space, we limit the GPU memory allocation choices for each worker to two cases: (a) the same as the non-parallelized setup, and (b) the minimal memory required to perform inference (proportional to the inverse of pipeline parallelism size). We call them full-memory workers and low-memory workers, respectively.
@@
Our algorithm has two stages. First, we enumerate all deployment choices with different pipeline parallelism sizes or GPU memory usage and predict the TTFT and worst-case TPOT for each choice. Second, we select the optimal choice that satisfies user SLOs with minimum resource usage. The maximum pipeline parallelism size is limited to as larger parallelism sizes yield little improvement.
@@
The TTFT and TPOT prediction takes historical information as the input. Formally, it includes the time cost of container creation and runtime initialization (), data transmission (), prefill (), and decoding (). The data transmission time is the latency of the TCP network between GPU servers. The prefill and decoding time costs are model-specific metrics obtained from the model’s previous executions. Assume the pipeline parallelism size for the current deployment choice is and there are full-memory workers , while the network and PCIe bandwidth of servers on which workers reside are and . Finally, we predict the TTFT using the following equation.
@@
where is the model size. Eq. 1 is made up of four parts: runtime initialization, model fetching, model loading, and prefilling. The model fetching and loading time is determined by the partitioned model size and speed of the slowest worker. During prefilling, a full-memory worker costs units of time, while a low-memory worker costs units.
@@
To satisfy TTFT requirement, we select GPU servers to minimize the model fetching and loading time for a given pipeline parallelism size and number of full-memory workers . Assume the GPU servers that can accommodate full-memory workers are . Apart from them, there are also servers that can accommodate low-memory workers, denoted as . Our selection strategy is to first allocate the top servers with minimum model fetching and loading time (i.e., the smallest ) from to full-memory workers, and then merge the remaining servers into the set . After that, we select the top servers from the merged set by the same strategy and allocate them to all low-memory workers.
Similar to the calculation of prefilling latency, we can predict the TPOT as follows.
Among all choices that satisfy both TTFT and TPOT SLOs, HydraServe prioritizes free GPUs during worker placement. When the cluster is not under heavy load, this approach improves the inference performance of models. Algorithm 1 outlines the resource allocation algorithm.
Input: time cost of container creation and runtime initialization , data transmission , prefill , and decoding ; model size ; GPU server network bandwidth and PCIe bandwidth ; user specified requirements and .
Output: pipeline parallelism size , #full-memory workers , and selected GPU servers .
@@ 4.2 Network-Contention-Aware Worker Placement
The second type of contention is the interference among cold-start workers on the same GPU server, which leads to unpredictable cold start performance. To solve this problem, we propose a network-contention-aware worker placement policy that places cold-start workers based on user SLOs.
Policy design. Initially, colocated workers share the network bandwidth with equal credits. Once we assign a worker to a GPU server, we record the model size it needs to fetch and the user-specified maximum TTFT. To place a new worker, we inspect each GPU server to check whether adding a cold-start worker to this server would lead to SLO violations for existing workers. If the check passes, we estimate and record the network bandwidth that the new worker can obtain. Note that PCIe bandwidth is much higher than network and PCIe switch is able to further isolate PCIe usage across tasks, so we do not take PCIe contention into account.
Specifically, we denote the bandwidth of a GPU server as . For each cold-start worker on the GPU server, we record the fetching deadline and estimate its pending model size . The fetching deadline comes from the prediction of TTFT (Eq. 1). When a new cold-start worker comes, the network bandwidth per work decreases, and we check whether all cold-start workers on the server are able to finish timely under the new bandwidth. Formally, we check whether the following condition is true for all workers.
where represents the number of existing workers on the server and represents the current time. This server accepts the new worker if all workers passed the check.
To estimate the pending model size, we record the time of last network bandwidth change ’. The start and completion of a cold start change the network bandwidth, informing the controller to adjust the pending model size. Here, we assume the current time is and the number of cold-start workers before change is . Then we adjust the pending model size upon each bandwidth change according to the following equation.
means that the worker has fetched the model ideally, thus we delete it from the cold-start worker list.
@@ 5 Worker-Level Overlapping
As described in §4.1, the TTFT of pipeline parallelism has diminishing marginal returns due to container initialization stages, including container creation, library loading and CUDA context initialization. In this section, we present the worker-level overlapping in HydraServe that carefully reorganizes the startup workflow to reduce worker initialization time. Basically, HydraServe leverages two optimizations to overlap initialization stages. First, we prefetch model weights on local GPU server as soon as container creation starts so that the container initialization stages are overlapped with model fetching. Second, we prioritize CUDA context initialization and develop a parameter manager to load model while initializing Python libraries so that the library loading stage is overlapped with model loading.
@@ 5.1 Model Prefetching
HydraServe launches a model prefetcher on each GPU server, which is responsible for proactively fetching models from remote storage. When a worker has been allocated to the server, the central controller informs the model prefetcher about model metadata. After that, the prefetcher starts to load the model weights from remote storage to a shared memory region. The cold-start worker will fetch parameters from shared memory in a streaming manner after runtime has initialized.
Model prefetching starts before container creation, so that the container creation and runtime initialization stages are overlapped. The worker performs model prefetching and loading in a pipelined fashion. In the shared memory region of a model, we use the first eight bytes to store the address that represents the end of currently fetched model weights. Model weights are represented using the SafeTensors format [64]. This format contains the metadata of all parameters at the beginning of the file, so that it is convenient for the worker to check whether a tensor has been fetched.
@@ 5.2 Parameter Manager
HydraServe leverages a parameter manager to load model parameters in the background. The parameter manager runs in an individual thread and is responsible for resolving tensor metadata, reading weights from the shared memory, and finally loading weights into the GPU. The whole procedure is zero-copy and pipelined. The parameter manager also takes advantage of the high parallelism of GPU cores and uses multiple CUDA streams to load models. The priorities of CUDA streams are determined according to whether the loading process is on the critical path or in the background.
@@
Figure 6(a) shows the optimized cold-start workflow after adopting model prefetching and parameter manager. Model fetching and runtime preparation runs in parallel, while library loading is overlapped with model loading. In this way, the overhead of other cold start stages is concealed behind model fetching, reinforcing the effectiveness of pipeline parallelism.
As HydraServe may launch workers in pipeline parallelism groups and later load the remaining part of model in background, Figure 6(b) also shows the cold-start workflow in this scenario. The model prefetcher downloads two parts of model sequentially. After the first part of model has been loaded, the worker starts to perform inference with other workers in the pipeline parallelism group. Meanwhile, the parameter manager loads the second part of model in background and finally transforms itself into a standalone worker.
TTFT prediction with worker-level overlapping. After applying worker-level optimizations, the TTFT is reduced by fine-grained overlapping. Based on additional historical information including the time cost of container creation (), CUDA context initialization () and library loading (), we change the TTFT prediction used in resource allocation (Eq. 1) into the following formula.
6 Inference-Level Consolidation
As HydraServe creates a group of workers in parallel when the model only needs one, we propose pipeline consolidation that merges pipeline parallelism groups into individual workers to obtain optimal performance. Specifically, after the parameter manager has loaded the requested model parameters in pipeline-parallel inference, we inform it to continue loading the remaining part of model in background. Once loading completes, the worker returns to the non-parallelized setup and serves subsequent requests with the whole model in local GPU, thereby achieving optimal performance. The second part of model is loaded in low-priority CUDA streams, so that the performance of inference task will not be affected.
@@ 6.1 Worker Scaling
HydraServe provides two scaling choices. First, after workers have loaded corresponding parts of models, we can perform scaling down by allowing only one of them to fetch the unloaded model parts in background. Once all parameters are loaded, we migrate all existing requests to that worker along with their key-value cache. Finally, the worker continues to generate tokens with whole model while other workers are terminated. Figure 4(c) demonstrates the process of scaling down. In this way, we produce the first token earlier by parallelized model fetching and finally obtain a worker with whole model, similar to standard cold starts.
The second scaling choice is scaling up, converting all cold-start workers into individual serving endpoints, as illustrated in Figure 4(d). This approach tackles load spikes, where the autoscaler aims to deploy multiple workers for the model simultaneously. By enabling the rapid creation of workers within pipeline parallelism groups, this design allows the system to achieve maximum throughput more quickly.
By default, HydraServe adopts the scaling down mechanism to reduce model fetching latency with minimal overhead. To promptly react to bursty loads and transition to the scaling up mechanism, we manage worker lifecycles with a sliding window strategy. For each model, the number of requests received in the previous window is recorded and used to predict the maximum number of requests likely to arrive in the next window. The required number of workers is then determined based on the current waiting queue length combined with the predicted maximum number of requests expected to arrive in the next window. In cold start, we create a pipeline parallelism group that is no smaller than required number of new workers and later transform the group into desired number of individual workers. To handle sudden request surges, multiple pipeline parallelism groups can be created as needed.